OpenStudy (osuxbucks15):

A triangle has side lengths of 34 in., 20 in., and 47 in. Classify it as acute, obtuse, or right right obtuse acute

OpenStudy (lgbasallote):

\[34^2 + 20^2 > 47^2 \rightarrow \text{acute}\] \[34^2 + 20 ^2 = 47^2 \rightarrow \text{right}\] \[34^2 + 20^2 < 47^2 \rightarrow \text{obtuse}\]

OpenStudy (lgbasallote):

try...and see which is which

OpenStudy (campbell_st):

use pythagoras' theorem 47^2 = 34^2 + 20^2 if a^2 = b^2 + c^2 ... right triangle a^2 < b^2 + c^2 ... obtuse a^2 > b^2 + c^2 .. acute

OpenStudy (osuxbucks15):

ok but how will i no which 1 it is?

OpenStudy (lgbasallote):

what is 34^2 + 20^2

OpenStudy (lgbasallote):

try it in your calculator

OpenStudy (osuxbucks15):

1156+400

OpenStudy (osuxbucks15):

1156 400 1556

OpenStudy (osuxbucks15):

1556

OpenStudy (lgbasallote):

right...and what's 47^2?

OpenStudy (osuxbucks15):

2209

OpenStudy (lgbasallote):

so which is bigger 2209 or 1156?

OpenStudy (osuxbucks15):

2209

OpenStudy (osuxbucks15):

is it acute

OpenStudy (lgbasallote):

so 34^2 + 20^2 < 47^2 which one is it?

OpenStudy (osuxbucks15):

acute

OpenStudy (lgbasallote):

are you sure?

OpenStudy (lgbasallote):

look at my first post again

OpenStudy (osuxbucks15):

?

OpenStudy (lgbasallote):

look at what i wrote for 34^2 + 20^2 < 47^2

OpenStudy (osuxbucks15):

is it obtuse

OpenStudy (lgbasallote):

yes

OpenStudy (osuxbucks15):

thanks

OpenStudy (lgbasallote):

welcome :D